Morgan Valley Inc. has current sales of $540,000. All of its sales are on credit and there is no default. It currently sells on terms of net 30 and has an accounts receivable balance of $50,000.

The company is considering a new credit policy with terms of net 60. Under the new policy, sales are expected to increase to $600,000, and accounts receivable are expected to increase to $100,000.

Part 1

What is the firm's days sales outstanding under the old policy?

Correct ✓

DSO=Accounts receivableDaily credit sales=Accounts receivableAnnual credit sales360DSO=Accounts receivableDaily credit sales=Accounts receivableAnnual credit sales360

= 50,000540,00036050,000540,000360

= 33.33 Correct ✓

Part 2

What is the firm's days sales outstanding under the old policy?
Answer #1

DSO = 360*Accounts Receivables/Credit sales

Under old policy, DSO = 360*50,000/540,000 = 33.33 days

Under new policy, DSO = 360*100,000/600,000 = 60 days